OpenStudy (anonymous):

write this equation in slope-intercept form of the line that passes through the given point and is perpendicular to the graph of the given equation (-3,2);x-2y=7

OpenStudy (nikato):

So the slope intercept formula is y=mx+b Can u somehow make x-2y=7 look like that^

OpenStudy (anonymous):

y=1/2x-7/2 perpendicular: flip the slope, for ex: 2/3 into -3/2 change the sign, from - to + and vice versa